OpenStudy (anonymous):

What is the period of the sinusoid given by y = -4sin(((2π)/3)x) ?

OpenStudy (anonymous):

when does \[\frac{ 2\pi }{ 3 }x= 2 \pi\]

OpenStudy (anonymous):

3 is the period!

OpenStudy (anonymous):

thwere you go

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

u r welcome